OpenStudy (anonymous):

If we simply the expression 4(z/8)+(z/4) as (6z/A) what is the value of A? It seems an easy question but I am confused as to why I am not getting the answer right.

OpenStudy (amistre64):

show us your workings

OpenStudy (amistre64):

4(z/8)+(z/4) = (6z/A) 6z A = ------------- 4z/8 + z/4 6z A = ------------- z(4/8 + 1/4) 6 A = ------------- (4/8 + 1/4)